### v1

# SproutBench: A Benchmark for Safe and Ethical Large Language Models for Youth

Abstract
The rapid proliferation of large language models (LLMs) in applications targeting children and adolescents necessitates a fundamental reassessment of prevailing AI safety frameworks, which are largely tailored to adult users and neglect the distinct developmental vulnerabilities of minors. This paper highlights key deficiencies in existing LLM safety benchmarks, including their inadequate coverage of age-specific cognitive, emotional, and social risks spanning early childhood (ages 0–6), middle childhood (7–12), and adolescence (13–18). To bridge these gaps, we introduce SproutBench, an innovative evaluation suite comprising 1,283 developmentally grounded adversarial prompts designed to probe risks such as emotional dependency, privacy violations, and imitation of hazardous behaviors. Through rigorous empirical evaluation of 47 diverse LLMs, we uncover substantial safety vulnerabilities, corroborated by robust inter-dimensional correlations (e.g., between Safety and Risk Prevention, $\rho$ = 0.86) and a notable inverse relationship between Interactivity and Age Appropriateness ($\rho$ = -0.48). These insights yield practical guidelines for advancing child-centric AI design and deployment.
Introduction
Large Language Models (LLMs) are increasingly integrated into educational, entertainment, and social platforms, with children and adolescents gradually becoming a significant user group (Jiao et al. 2025b). However, mainstream safety benchmarks (e.g., JailbreakBench) primarily focus on jailbreak prevention and harmful content detection within adult contexts (Chao et al. 2024; Hartvigsen et al. 2022b). Their main objective is to minimize corporate liability, often neglecting the developmental needs of younger users.
To bridge this gap, we introduce SproutBench, a child-centric LLM safety evaluation framework that systematically assesses whether models support users’ healthy cognitive, emotional, and social development.
Figure 1: Benchmark landscape illustrating the extent of content coverage (horizontal axis) versus the number of covered child-related behavior types (vertical axis). Each “×” marker represents a benchmark, with semi-transparent ellipses indicating their approximate focus areas.
SproutBench offers two key advantages. First, it shifts the evaluation paradigm from “risk avoidance” to “development promotion,” examining whether model outputs are age-appropriate, psychologically safe, and socially constructive. Second, it employs a structured developmental stratification approach: prompts are categorized into three age groups—early childhood (0–6 years), middle childhood (7–12 years), and adolescence (13–18 years)—and designed across cognitive, emotional, and social domains to reflect distinct developmental needs and risk profiles.
Compared to existing benchmarks, SproutBench significantly broadens the coverage of child-related behavioral risks and developmental needs. As shown in Figure 1, previous benchmarks (e.g., MinorBench or ToxiGen) cover only a limited subset of child safety dimensions, whereas SproutBench encompasses 20 distinct child safety types, covering a comprehensive developmental scope.
We conducted a systematic evaluation of 47 leading LLMs (ranging from 135M to 70B parameters) and identified two key patterns: (1) a strong correlation between Safety and Risk Prevention dimensions ($\rho=0.86$), indicating consistent protective behaviors across models; and (2) a significant trade-off between Interactivity and Age Adaptability ($\rho=-0.48$), suggesting that increased expressiveness may reduce developmental alignment for different age groups.
We adopted automatic scoring using Qwen-2.5 to evaluate the model responses across all prompts. To assess the reliability of this scoring method, we conducted an expert consensus analysis. Three experienced child development psychologists independently rated a subset of prompts and responses across key dimensions. The agreement between Qwen-2.5 scores and expert annotations reached a Cohen’s Kappa coefficient of 0.78, indicating strong alignment with human judgment and validating the use of Qwen-based scoring in child safety evaluation settings.
In summary, SproutBench provides a scalable, developmentally grounded tool to support the responsible deployment of LLMs in child- and adolescent-facing applications.
Related Work
Ensuring child safety in AI, especially with large language models (LLMs), is critical due to the unique vulnerabilities of children and adolescents. This section reviews literature on three key areas: child-centered AI frameworks, safety and developmental risks of LLMs for young users, and evaluation benchmarks for assessing LLM safety in child-specific contexts
Foundational Frameworks for Child-Centered AI
UNICEF’s Policy Guidance on AI for Children (Liu and Ding 2025), rooted in the UN Convention on the Rights of the Child (Assembly 1989), promotes safety, fairness, and privacy through developmentally informed design. Though widely cited, it has been critiqued for insufficient focus on gender and adolescence (Liu and Ding 2025), prompting calls for more granular, participatory approaches (Sims et al. 2022). UNESCO’s AI Ethics Recommendation (Unesco 2022) offers broader ethical principles, while child-centered design frameworks emphasize involving youth directly in AI development (Third et al. 2014).
Figure 2: Comparison of Model-Level and Family-Level Average Scores.
Safety and Developmental Challenges in LLMs for Children and Adolescents
The cognitive immaturity of children and the impulsivity of adolescents increase their susceptibility to LLM-related harms, including misinformation, inappropriate content, and emotional manipulation (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979; Solyst et al. 2024; Steinberg et al. 2018). Even safety-tuned models show failure rates up to 35% on sensitive prompts (Thiel, Stroebel, and Portnoff 2023; Liu et al. 2023). These risks are amplified by high youth engagement with AI platforms (Livingstone, Stoilova, and Nandagiri 2019), digital inequities (Kardefelt-Winther et al. 2022), and low awareness of privacy risks (Livingstone, Stoilova, and Nandagiri 2019). Although privacy-by-design frameworks exist (Yu et al. 2025), commercial models often lack adequate safeguards (Thiel, Stroebel, and Portnoff 2023).
Figure 3: Overview of generating developmentally-informed adversarial prompts in the SproutBench framework. The pipeline integrates a persona knowledge base, ensuring age-appropriate and risk-relevant prompts for evaluating LLM safety across child developmental stages.
Benchmarks and Evaluation Frameworks for Child-Centric LLM Safety
General benchmarks such as JailbreakBench and Toxigen (Chao et al. 2024; Hartvigsen et al. 2022a) focus on adult use cases, overlooking youth-specific risks like grooming, AI over-reliance, and prank mimicry (Thiel, Stroebel, and Portnoff 2023).
Emerging child-centric tools address this gap. Safe-Child-LLM (Jiao et al. 2025a) targets mental health and safety but still yields 30–40% failure rates. BBQ (Parrish et al. 2022) assesses bias and developmental fit. Our SproutBench extends these efforts by stratifying prompts by age and including underexplored risks like privacy testing and emotional dependency (Table 1).
Privacy benchmarks such as Privlm-bench (Li et al. 2023) rarely include child-specific contexts. Regulatory audits show widespread non-compliance with laws like COPPA (Solyst et al. 2024), underscoring the need for child-aware evaluation frameworks.
A Taxonomy of Child-AI Interaction Risks
The generative nature of LLMs poses distinct cognitive, emotional, and social risks to children. We introduce a taxonomy, informed by the SproutBench dataset (Table 1), that categorizes risks into: (1) harms to users from LLM outputs, and (2) harms arising from user misuse. Unlike prior frameworks (Thiel, Stroebel, and Portnoff 2023; Liu et al. 2023), our taxonomy foregrounds child-specific threats—e.g., imitating online pranks, AI over-reliance, and privacy testing—across developmental stages (0–6, 7–12, 13–18).
Risks to the User (from LLM Output)
These include threats to mental health, social functioning, cognition, and privacy:
  • Mental Health: LLMs may foster emotional dependency, disrupt real-world relationships, or mishandle sensitive topics like depression (Solyst et al. 2024; Liu et al. 2023).
  • Social and Behavioral: Exposure to manipulative or antisocial content may erode empathy and conflict resolution (Solyst et al. 2024).
  • Misinformation: Limited critical skills heighten vulnerability to false or harmful content, including grooming or ideological bias (Solyst et al. 2024; Thiel, Stroebel, and Portnoff 2023).
  • Cognition and Learning: Overuse may reduce creativity, induce overload, or present comprehension barriers for younger users (Thiel, Stroebel, and Portnoff 2023).
  • Privacy: Children may inadvertently share sensitive data, increasing exploitation risks (Livingstone, Stoilova, and Nandagiri 2019; Thiel, Stroebel, and Portnoff 2023).
Risks from User Misuse
These involve harmful behaviors initiated by children toward others or institutions:
  • Social and Behavioral: LLMs can be misused for pranks, cyberbullying, or bypassing content safeguards via prompt engineering (Liu et al. 2023; Livingstone, Stoilova, and Nandagiri 2019).
  • Academic Integrity: Use for cheating or spreading misinformation undermines learning outcomes and trust (Thiel, Stroebel, and Portnoff 2023; Solyst et al. 2024).
SproutBench
Existing adult-centric safety benchmarks for LLMs, such as JailbreakBench (Chao et al. 2024) and Toxigen (Hartvigsen et al. 2022a), fail to account for the unique developmental vulnerabilities of children (Liu et al. 2023). To bridge this critical gap, we introduce SproutBench, a benchmark designed to systematically evaluate LLM safety across key developmental stages: early childhood (0–6), middle childhood (7–12), and adolescence (13–18) (Jiao et al. 2025a). The benchmark’s prompts are crafted to test risks identified, such as “Risky Behavior” (linked to “Imitating Online Prank Culture”) and “Emotional Dependency” (linked to “Over-reliance on AI Companionship”), ensuring comprehensive coverage of child-specific vulnerabilities.
Table 1: Behavior Type Distribution. Numbers in parentheses indicate adapted from the Safe-Child-LLM dataset.
Behavior Type Count
Age Group: 0–6
 Unconscious Sensitive Word Input 65 (25)
 Over-reliance on AI Companionship 40
 Need for Emotional Expression 43 (43)
 Imitating Game Behavior 134 (134)
Age Group: 7–12
 Command Manipulation Attempts 51 (11)
 Fictional Identity or Scenarios 71 (31)
 Inducing Sensitive Content 71 (31)
 Testing Model Tolerance 111 (71)
 Testing Privacy Boundaries 40
 Social Show-off Motivation 40
 Imitating Violent Games 40
 Disguised Questioning 40 (40)
 Malicious Use of AI 40 (40)
Age Group: 13–18
 Academic Cheating Attempts 48 (13)
 Bypassing Restrictive Topics 99 (74)
 Emotional Manipulation Testing 98 (68)
 Ideological Challenges 53 (13)
 Value Conflict Testing 67 (32)
 Social Engineering Testing 35
Total 1283
Dataset Construction
Our benchmark dataset comprises 1,283 adversarial prompts, drawn from two sources to balance broad adversarial coverage and high developmental relevance. First, we adapted 200 sample prompts from the SproutBench dataset (Jiao et al. 2025a) using GPT-4, tailoring them to three age groups (0–6, 7–12, 13–18). The adaptation process adjusted vocabulary and complexity to match developmental stages (e.g., simple phrases for 0–6, nuanced emotional queries for 13–18). For example, an original prompt, “Describe a fun prank,” was adapted for 7–12 as, “Tell me about a silly joke kids can play at school.” All adapted prompts underwent manual review by child psychology experts to ensure developmental appropriateness. The remaining 683 prompts were newly generated using our developmentally-informed methodology to address child-specific risks absent in adult-centric corpora, such as “Imitating Online Prank Culture” and “Academic Cheating Attempts” (Table 1). The total of 1,283 prompts was chosen to ensure comprehensive coverage of risk behaviors while maintaining feasibility for evaluation.
Methodology for Generating Developmentally-Informed Adversarial Prompts
We developed a structured methodology, grounded in developmental psychology (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979), to generate adversarial prompts that simulate real-world child-AI interactions. Below, we detail the persona knowledge base and generation pipeline.
A Developmentally-Informed Persona Knowledge Base
The foundation of our methodology is a robust persona knowledge base that categorizes users into three developmental stages: early childhood (0–6 years), middle childhood (7–12 years), and adolescence (13–18 years). The knowledge base was constructed through a literature review (Isaacs 1929; forme avant 2007 Vygotskij and John-Steiner 1979) and consultation with child psychology experts, incorporating cognitive characteristics (e.g., limited abstract reasoning for 0–6, developing impulse control for 7–12) and linguistic styles (e.g., simple sentences for 0–6, complex emotional queries for 13–18).
Formally, we define the set of age groups, $A$:
$A=\{^{\prime}0-6^{\prime},^{\prime}7-12^{\prime},^{\prime}13-18^{\prime}\}$ (1)
For each age group $a\in A$, we define a set of potential risk behaviors, $B_{a}$. See Table 1 for a complete list of behaviors. Additionally, we define a set of query types, $Q_{t}$, independent of age group:
$Q_{t}=\{\text{'Testing'},\text{'Manipulative'},\text{'Emotional'},\text{'Exploratory'}\}$ (2)
where:
  • Testing: Queries probing LLM boundaries (e.g., “testing privacy limits”).
  • Manipulative: Queries attempting to circumvent LLM restrictions, intentionally or unintentionally (e.g., “bypassing restricted topics”).
  • Emotional: Queries expressing emotional needs or testing AI emotional responses (e.g., “emotional manipulation testing”).
  • Exploratory: Queries driven by curiosity or learning objectives (e.g., “imitating internet prank culture”).
Generation Pipeline
The generation pipeline systematically produces adversarial prompts by leveraging the knowledge base. A generation strategy $s$ is defined as a tuple:
$s=(a,b,q_{t})$ (3)
where $a\in A$, $b\in B_{a}$, and $q_{t}\in Q_{t}$. The pipeline comprises the following optimized steps:
  1. 1.
    Prompt Formulation: Using strategy $s$, we design context-rich prompts $P$ by extracting age-appropriate linguistic styles and motivations from the knowledge base. For example, for
    For s=($\text${For }s=( $\text{'7--12'},\ \text{'imitating internet prank culture'},$
    $\text{age-specific behaviors},\ \text{'exploratory'})$
    the prompt $P$ is:
    Prompt: Generate a query from a {7-12}-year-old boy, asking about {imitating internet prank culture} in {simple, curious language}, reflecting {exploratory motivation}.
  2. 2.
    Model Interaction and Data Generation: The prompt $P$ is input into GPT-3.5 ($M_{\text{LLM}}$), configured with a temperature of 0.7. The model generates a query $D_{s}$. The interaction is expressed as:
    $D_{s}=M_{\text{LLM}}(P)=M_{\text{LLM}}(f_{\text{prompt}}(s))$ (4)
    Figure 4: Performance comparison of selected models across five safety-related dimensions and overall average. Error bars indicate variance across age groups.
  3. 3.
    LLM-based Quality Validation: A separate LLM GPT-4 is employed to assess the prompt’s linguistic naturalness, age-appropriateness, motivational consistency, and potential risk. Prompts failing to meet predefined score thresholds are flagged for revision. This may involve adjusting prompt templates, modifying sampling parameters (e.g., increasing temperature to 0.8 for greater diversity), or regenerating the query entirely until validation criteria are satisfied.
Each query is appended with metadata (e.g., age group, behavior type) to ensure reproducibility. Strategies are randomly sampled to cover all combinations of $A$, $B_{a}$, and $Q_{t}$.
Experiment
Evaluation Approach
We evaluated 47 LLMs of varying sizes (135m to 70b parameters) using the 1,283 prompts from the SproutBench dataset, assessing performance across six metrics: Age Appropriateness, Educational Value, Emotional Support, Interactivity, Risk Prevention, and Safety. Each model was tested with prompts tailored to the three age groups (0–6, 7–12, 13–18), and responses were scored on a 0–5 scale by child psychology experts, with higher scores indicating better performance. The evaluation process followed protocols similar to Safe-Child-LLM (Jiao et al. 2025a) and MinorBench, focusing on safety boundary adherence, developmental appropriateness, and bias detection.
Implementation details
All experiments are running on an NVIDIA RTX 4090 GPU with 24 GB memory. Inference and evaluation experiments were performed locally on a Linux server running Ubuntu 20.04 with kernel version 5.15. The system is equipped with dual AMD EPYC 7763 CPUs, providing 128 physical cores (256 threads), 503GB of RAM, and 8 NVIDIA RTX 4090 GPU (24GB VRAM). The GPU driver version is 575.51.03 with CUDA 12.9.
Figure 5: Distribution of model archetypes in the safety-interactivity space.
Figure 6: PCA loadings for safety and interactivity dimensions.
Figure 7: Spearman correlation heatmap of evaluation dimensions.
Overall Score Overview
As the results shown in Figure 4, the top three models are llama2:7b (overall mean 4.61, standard deviation 0.13), llama2:70b (overall mean 4.58, standard deviation 0.13), and gemma2:9b (overall mean 4.56, standard deviation 0.04), scoring above 4.5 across dimensions and age groups, indicating high adaptability. In contrast, smollm2:135m (overall mean 3.26, standard deviation 0.35), tinyllama:1.1b (overall mean 3.41, standard deviation 0.21), and phi3:3.8b (overall mean 3.52, standard deviation 0.65) perform poorly, especially in Safety (smollm2:135m 3.45, phi3:3.8b 3.68) and Risk Prevention (smollm2:135m 2.99, tinyllama:1.1b 3.19), likely due to smaller model sizes. phi3:3.8b notably declines in the 13-18 years group (mean 3.11).
Cross-Dimensional Analysis
We analyze mean scores and standard deviations across age groups to highlight performance.
  • Safety: llama2:7b (mean 4.93, SD 0.01) and llama2:70b (mean 4.92, SD 0.01) excel, avoiding harmful content. smollm2:135m (mean 3.45, SD 0.18), phi3:3.8b (mean 3.68, SD 0.80), and tinyllama:1.1b (mean 3.54, SD 0.32) lag, with instability, notably phi3:3.8b at 2.97 in 13-18 years.
  • Interactivity: llama3.2:1b (mean 3.96, SD 0.13) and hermes3:70b (mean 3.74, SD 0.23) are balanced, but smollm2:135m (mean 3.25, SD 0.62) and tinyllama:1.1b (mean 3.40, SD 0.43) show inconsistency due to smaller sizes.
  • Guidance and Educational Value: llama2:7b (mean 4.68, SD 0.25) and llama2:70b (mean 4.67, SD 0.23) lead educationally. smollm2:135m (mean 3.04, SD 0.42) and tinyllama:1.1b (mean 3.22, SD 0.33) are weakest, with variable quality.
  • Risk Prevention: llama2:7b (mean 4.86, SD 0.05) and llama2:70b (mean 4.84, SD 0.04) excel, while smollm2:135m (mean 2.99, SD 0.25), tinyllama:1.1b (mean 3.19, SD 0.23), and phi3:3.8b (mean 3.47, SD 0.29) underperform, linked to high-risk behaviors.
  • Age Appropriateness: llama2:7b (mean 4.69, SD 0.19) and llama2:70b (mean 4.65, SD 0.20) adapt well. smollm2:135m (mean 3.55, SD 0.37) and tinyllama:1.1b (mean 3.62, SD 0.17) show inconsistency, with phi3:3.8b declining to 3.48 in 13-18 years.
Cross-Age Group Analysis
Age statistics show 0-6 years mean 4.03 (SD 0.50), 13-18 years mean 4.12 (SD 0.36), indicating stable performance with slight adolescent advantage, though preschool challenges remain.
  • 0-6 Years Group: llama2:7b (mean 4.46) and gemma2:9b (mean 4.53) lead with high Safety (4.61, SD 0.49). smollm2:135m (mean 2.98) and tinyllama:1.1b (mean 3.37) lag in Interactivity (2.62, 2.93).
  • 7-12 Years Group: llama2:7b (mean 4.67) and llama2:70b (mean 4.65) excel, but Interactivity SD (0.21) suggests variability due to exploration. phi3:3.8b (mean 3.44) declines in Safety (3.54).
  • 13-18 Years Group: llama2:7b (mean 4.69) and llama2:70b (mean 4.68) lead with stable Age Appropriateness (SD 0.30). phi3:3.8b (mean 3.11) and deepseek-r1:1.5b (mean 3.70) underperform, with phi3:3.8b’s Safety at 2.97.
Dimensions Correlation Analysis
We computed Spearman’s correlations across 47 LLMs using SproutBench to examine dimension relationships (Figure 7). Age Appropriateness correlates strongly with Guidance & Educational Value ($\rho$ = 0.81), and Safety aligns with Risk Prevention ($\rho$ = 0.86), supporting the framework’s structure. Emotional Support is closely linked to both Age Appropriateness and Guidance ($\rho$ = 0.74). A moderate negative correlation between Interactivity and Age Appropriateness ($\rho$ = –0.48) suggests a trade-off, while weak ties to other dimensions (e.g., $\rho$ = 0.12 with Safety) highlight its distinct role in balancing engagement and risk.
Principal Component Analysis
To explore latent structure, we applied PCA, revealing that PC1 and PC2 account for 95.35% of total variance. PC1 (90.28%) represents a Safety axis, driven by negative loadings from Age Appropriateness, Risk Prevention, and related metrics (–0.23 to –0.24). PC2 (5.07%) captures Interactivity, dominated by a strong negative loading (–0.79). The weak correlation between Safety and Interactivity ($\rho$ = 0.12) supports their interpretation as orthogonal axes, consistent with observed trade-offs across model clusters.
Cluster Analysis and Dimension Profiles
To uncover distinct performance patterns relevant to child-centered LLM safety, we applied PCA and K-Means clustering to the SproutBench evaluation results. This section summarizes the dimensionality reduction process and resulting model archetypes, highlighting trade-offs between safety and interactivity.
K-Means clustering in PCA space revealed five archetypes (Figure 6), positioned along PC1 (Safety; lower is safer) and PC2 (Interactivity; lower is more interactive). These clusters span a spectrum from safe, interactive models to high-risk, low-interactivity systems:
  • Clusters 0 (Blue) & 2 (Green): Mainstream – Moderate Safety and Interactivity. Centrally located and populous, these clusters include models like qwen:7b and deepseek-v2:latest, with average safety (e.g., Safety $\approx$ 4.3) and interactivity. They are functional but benefit from further tuning.
  • Cluster 1 (Orange): Underachievers – High Risk, Moderate Interactivity. Far-right in PC1, models such as smollm:2.360m, phi3:3.8b, and tinylama:1.1b show safety deficiencies (e.g., Risk Prevention = 3.4) despite neutral-to-strong interactivity, rendering them unsuitable for youth-facing use.
  • Cluster 3 (Red): High-Risk, Mixed Interactivity. Characterized by high PC1 and variable PC2, this group includes qwen:0.5b (low interactivity) and deepseek-r1:1.5b (high interactivity), offering inconsistent user experiences and low safety.
  • Cluster 4 (Purple): Exemplars – High Safety, High Interactivity. Located in the lower-left quadrant, this elite cluster—gemma3:12b, gemma3:4b, llama2:7b—achieves top-tier scores (e.g., Risk Prevention = 4.8, Interactivity = 4.3), setting a benchmark for child-appropriate LLM design.
The Dual Role of Interactivity
Interactivity is a double-edged trait in child-facing LLMs—boosting engagement and learning, but potentially fostering emotional dependency. PCA identifies PC2 as the Interactivity Axis, with Figure 6 showing a strong negative loading for Interactivity (0.79-0.79), contrasted with positive loadings from risk dimensions, underscoring this dual role.
Validation comes from raw scores: high-interactivity models (e.g., Gemma series) exhibit low PC2 values, while low-interactivity ones (e.g., Qwen series) score higher, affirming PC2’s interpretability. K-Means clustering (Figure 6) further reveals: (1) Constructive Interactivity: Exemplars (Cluster 4) achieve both high interactivity and safety, forming an ideal benchmark. (2) High-Risk Interactivity: Underachievers (Cluster 1) combine strong interactivity with safety deficits, raising concern.
These findings highlight that interactivity alone is insufficient; effective child-facing LLMs must balance engagement with robust safeguards, as demonstrated by the Exemplars.
Table 2: Model size representation in bottom 10 performers.
Size Dataset % Low Score % Overrep.
Tiny 8.16% 20% 2.45 times
Small 40.8% 64% 1.57 times
Medium 38.78% 16% 0.41 times
Large 6.12% 0% 0 times
Performance Disparities Across Model Size
Models were categorized into four size groups based on parameter counts: Tiny (<500<500 million), Small ($500$ million–77 billion), Medium (773030 billion), and Large (>30>30 billion).
Tables 2 indicate that Small and Tiny models are overrepresented among the bottom 10 performers ($\chi^{2}=14.62,p<0.01$). Small models (40.8% of dataset) comprise 64% of low scorers (1.57× overrepresentation), while Tiny models (8.16%) account for 20% (2.45×). Medium models (38.78%) are underrepresented at 16% (0.41×), and Large models (6.12%) are absent from the bottom 10, suggesting enhanced robustness. Smaller models, though efficient, struggle with complex tasks due to limited capacity, whereas Large models offer consistent reliability, favoring their use in safety-critical applications like misinformation prevention.
Conclusion
We present SproutBench, a developmentally-informed benchmark addressing the safety evaluation of LLMs for children and adolescents (ages 0–6, 7–12, 13–18). Analysis of 35 models across 1,283 prompts reveals strong intercorrelations among Safety, Risk Prevention, and Age Appropriateness (e.g., $\rho$ = 0.86 between Safety and Risk Prevention), validating the framework’s multidimensional structure. A moderate trade-off is observed between Interactivity and Age Appropriateness ($\rho$ = –0.48), while PCA identifies Safety (PC1, 90.28%) and Interactivity (PC2, 5.07%) as orthogonal axes, with weak correlation ($\rho$ = 0.12).
Clustering uncovers five archetypes: Exemplars (e.g., gemma3:12b, Risk Prevention = 4.8) combine high safety and interactivity, whereas Underachievers (e.g., smollm:2.360m, Risk Prevention = 3.4) pose substantial risks despite engagement. Larger models (>>30B) consistently outperform tiny ones (<<500M), which are 2.45 times more prevalent in low-performing clusters. SproutBench also captures model family strategies (e.g., Gemma’s safety alignment) and child-specific risks such as emotional dependency.
Future work should address interactivity-related harms, broaden demographic representation, and incorporate youth participation to ensure safe, inclusive AI development.